OpenStudy (anonymous):

In order for a button to fit through its buttonhole, the hole needs to be the size of the button's diameter. What size buttonhole is needed for a button with a circumference of 9.42 centimeters?

OpenStudy (cwrw238):

circumference = diameter * pi 9.42 = d * pi d = 9.42 / pi

OpenStudy (anonymous):

3

OpenStudy (anonymous):

thank you guys

OpenStudy (cwrw238):

yw